A marketing research company desires to know the mean consumption of meat per week among males over age 27. They believe that the meat consumption has a mean of 3.2 pounds, and want to construct a 90% confidence interval with a maximum error of 0.07 pounds. Assuming a standard deviation of 0.9 pounds, what is the minimum number of males over age 27 they must include in their sample? Round your answer up to the next integer.
Answer #1

Solution:

Given ,

= 0.9 ..Population SD

E = 0.07 Margin of error

c = 90% = 0.90 ...confidence level

Find sample size required.

c = 0.90

= 1 - c = 1 - 0.90 = 0.10

/2 = 0.05

= 1.645 (using z table)

Now, sample size (n) is given by,

=  {(1.645* 0.9)/ 0.07}2

=  447.3225

= 448 ..(round to the next whole number)

Answer : Required Sample size is n = 448